Find the inverse Laplace transform of each of the following functions. In several of these problems, it will be helpful to write the functions 2 (as b)(cs2+ds +e) in the simpler form As B as2 + bs + c ds2+ es + f Cs+ D CsD and p2(s) = as + b T cs2 + ds + e 2
Answer #1

Solution: Let \phi \left ( s \right )= \frac{s}{\left ( s+a \right )^{2}+b^{2}}

Apply Inverse laplace on both side

L^{-1}\phi \left ( s \right )=L^{-1}\left [ \frac{s}{\left ( s+a \right )^{2}+b^{2}} \right ]

     =L^{-1}\left [ \frac{s+a-a}{\left ( s+a \right )^{2}+b^{2}} \right ]

     =L^{-1}\left [ \frac{s+a}{\left ( s+a \right )^{2}+b^{2}} -\frac{a}{\left ( s+a \right )^{2}+b^{2}}\right ]

     =L^{-1}\left [ \frac{s+a}{\left ( s+a \right )^{2}+b^{2}} \right ] -aL^{-1}\left [ \frac{1}{\left ( s+a \right )^{2}+b^{2}} \right ]\left ( \because \mathrm{Linearity \: property }\right )

=e^{-at}L^{-1}\left [ \frac{s}{ s^{2}+b^{2}} \right ] -ae^{-at}L^{-1}\left [ \frac{1}{ s^{2}+b^{2}} \right ]\left ( \because \mathrm{first \:shifting\: \: property }\right )

=e^{-at}cosbt -\frac{a}{b}e^{-at}sinbt